# Multi-Source Music Generation with Latent Diffusion

### v1

@@ IV-B FAD
For the task of total generation, we evaluate the generation of all the instrumental sources (piano, drums, bass, guitar) and the summed mixture. The FAD figure is shown in Fig. 2. First, for single-source generation, ISLDM in general has the lowest FAD, showing the best generation performance. This is reasonable because each independent diffusion model only needs to model one single source. Then our proposed MSLDM and MSLDM-Large also show promising results, beating the MSDM baseline by a large margin. The drums generated by MSLDM-Large even have FAD much smaller than the ISLDM model. For mixture generation, the MixLDM has the worst performance. This confirms our assumption that modeling the mixture directly is a much harder task than modeling sources. Then ISLDM, MSLDM, MSLDM-Large all have FAD much smaller than the MSDM, showing the effect of latent diffusion in modeling more realistic music mixtures. Then we can observe that MSLDM and MSLDM-Large have lower FAD than ISLDM, showing that our model is successfully generating sources with harmony. Note that when generating each single source, MSLDM performs worse than ISLDM for all instruments, but when sources are added together to form mixtures, MSLDM has lower FAD, showing that inter-source harmony is modeled.
For the task of partial generation, similar to MSDLM, we use the Sub-FAD as the metric, which is the FAD on the final mixture (partially generated sources + sources given). The reference samples for the FAD are then the ground truth mixtures. For the dataset used for FAD calculation, we again sample 200 12-second chunk music from slakh2100 test set, but we make sure that all the chunks contain four instruments. The Sub-FAD result is shown in Table II, where the results are shown for all models and all partial generation setups. The ISLDM model in this case is only partially generating the sources with the specific instrument model, without using any of the given sources. Then from the result table, overall, MSLDM-Large and MSLDM show much better performance than ISLDM and MSDM, with FADs smaller than 1. Interestingly, the ISLDM’s overall FAD is lower than MSDM, which means even ISLDM cannot generate consistent sources based on given sources, it’s able to model single-source melody much better than MSDM. For all the detailed setups, we can see that MSLDM is consistently better than MSDM and ISLDM, except for the guitar.
@@ IV-C Subjective Listening Test
To complement the objective metrics, we also design subjective tests. We follow the exact test design in MSDM [28] for total and partial generation. For the total generation subjective test, each model generates 10 chunks of 12-second music mixture samples. Then for each sample in the test, the participant is asked to rate the quality and coherence of each sample, with a score from 1 to 10 (higher the better). Quality corresponds to how realistic the music is and coherence corresponds to how coherent the sources inside the mixture are. For partial generation, for each model, we randomly sample 10 samples in the test set and then for each sample, we randomly sample the sources to generate. We provide the given conditional music, partially generated music, and the synthesized mixture, and then ask the participant to rate 1-10 for quality and density. The quality here means how coherent the generated sources are when mixed with the given sources. The density corresponds to how much the generated sources are in the 12-second chunk.
The subjective score is shown in Figure 3. For total generation, MSLDM and MSLDM-Large lead other baselines by a large margin in both quality and coherence. Also, MSLDM-Large has a much higher score than MSLDM, showing that large model size results in much better performance. The ISLDM’s quality is better than MSDM, but its coherence is the worst because all the sources are generated independently. Then MixLDM has the worst quality, showing the difficulty in directly modeling the music mixture, but its coherence is similar to MSDM. In general, users find that our MSLDM is much better than MSLDM, ISDM, MixLDM in both quality and coherence.
For partial generation, we can see that MSLDM and MSLDM-Large have the highest quality, showing the best performance in generating consistent sources that match the given ones. Again, ISLDM shows the worst quality because sources are independent. However, for density, ISLDM has the highest score showing independent source generation models can generate very active sources. Then MSLDM and MSLDM-Large also have a density score larger than 7 on average, showing the ability to generate active companions. However, MSDM shows the worst performance in density because it often generates small and naive music segments or empty sources.
Overall, both the objective metrics and subjective listening test show the MSLDM is much better than MSDM, ISLDM, MixLDM in terms of generating multi-source consistent music pieces. Compared with MSDM, MSLDM is better in terms of generating longer, more melodic, and more harmonious music. Compared with MixLDM, we showed that modeling sources is much easier than directly modeling music mixtures. Compared with ISLDM, we showed that our model is able to model inter-source dependency or harmony between different musical instruments.
@@ V Conclusion and Future Work
In this paper, we propose a multi-source diffusion model to jointly model the generation of multiple instruments together. Both objective and subjective metrics show much better results in the tasks of total and partial generation, implying MSLDM is able to model melody and source relations much better. Future research would include how to use MSLDM for weakly-supervised music separation and how to generalize our model to more instruments.

### v3

@@ IV-B Fréchet Audio Distance
For the task of total generation, we evaluate the generation of all the instrumental sources (piano, drums, bass, guitar) and the summed mixture (note: MixLDM can only be reported for the mixture). The FAD results are reported in Fig. 2. First, for single-source generation, ISLDM in general has the lowest FAD, showing the best generation performance. This is reasonable because each independent diffusion model only needs to model one single source. Then our proposed MSLDM and MSLDM-Large also show promising results, beating the MSDM baseline by a large margin. The drums generated by MSLDM-Large even achieve lower FAD than ISLDM. For mixture generation, MixLDM exhibits the worst performance. This confirms our belief that modeling the mixture directly is much harder than modeling sources. Observe that ISLDM, MSLDM, MSLDM-Large all outperform MSDM, demonstrating the efficacy of latent diffusion in modeling the harmony in realistic music mixtures. In fact, MSLDM and MSLDM-Large demonstrate lower FAD than ISLDM, showing that our model successfully generates sources in harmony. Note that when generating each single source, MSLDM performs worse than ISLDM for all instruments, but when sources are added together to form mixtures, MSLDM has lower FAD, showing that it is able to model the inter-source harmony.
For partial generation, we use the sub-FAD as the metric, similar to MSDM. sub-FAD calculates the FAD on a reference set and an evaluation set, where the reference set is the real music mixture, and the evaluated set is the mixture formed by mixing partially generated samples and originally given samples. For the dataset used for sub-FAD calculation, we again sample 200 12-second music segments from the slakh2100 test set, but we make sure that all the segments contain four instruments. The sub-FAD result is shown in Table II, where the results are shown for all models and all partial generation setups. The ISLDM model in this case is only generating the target sources using source-independent models. Overall, MSLDM-Large and MSLDM show much better performance than ISLDM and MSDM, with sub-FADs smaller than 1. Interestingly, the ISLDM’s overall sub-FAD is lower than MSDM, implying that even though ISLDM cannot generate mutually coherent sources based on given sources, it is able to model single-source melody much better than MSDM. Across all the detailed setups, we see that MSLDM is consistently better than MSDM and ISLDM, except for the guitar.
@@ IV-C Subjective Listening Test
To complement the objective metrics, we also design subjective listening tests with human evaluation. We follow the exact test design in MSDM [29] for total and partial generation.
Total Generation: Each model generates 10 segments of 12-second music mixture samples. Then for each sample in the test, the participant is asked to rate the ‘quality’ and ‘coherence’, with a score from 1 to 10 (higher the better). ‘Quality’ corresponds to how realistic the music is (considering white noise is the least realistic) and ‘coherence’ corresponds to how mutually coherent the sources inside the mixture are. The evaluation scores are shown in the upper half of Figure 3. MSLDM and MSLDM-Large lead other baselines by a large margin in both quality and coherence. Also, MSLDM-Large exhibits a higher score than MSLDM, showing that large model size results in better performance. ISLDM’s quality is better than MSDM, but its coherence is the worst because all the sources are generated independently. MixLDM has the worst quality and coherence, showing the difficulty in directly modeling the music mixture. In general, users find that our MSLDM model is consistently better than MSLDM, ISDM, MixLDM in both quality and coherence.
Partial Generation: We randomly sample 10 samples in the test set. Then, for each sample, the target source types (e.g., want to generate piano and drums given the two other instruments) are randomly sampled. Finally, the models are used to sample the specified source type for each data sample. We give the participants three music segments (the music to condition on, the partially generated music, and the synthesized mixture), and then ask them to rate 1-10 for ‘coherence’ and ‘density’. The ‘density’ corresponds to how dense (or sparse) the generated instruments are, in the 12-second segment (10 means all target instruments are generated for the whole segment). The results are shown in the lower half of Fig. 3. We observe that MSLDM and MSLDM-Large produce the highest coherence score, achieving the best performance in generating mutually-coherent sources that match the given ones. Again, ISLDM shows the worst coherence because sources are independent. For density, MSLDM, MSLDM-Large, and ISLDM all have high scores, showing the ability to generate rich companions. However, MSDM shows the worst performance in density because it often generates small and naive music segments, or even empty sources at times.
Overall, both the objective metrics and subjective listening test show the MSLDM outperforms MSDM, ISLDM, MixLDM in terms of generating multi-source consistent music pieces. Compared with MSDM, MSLDM is better in terms of generating denser, more melodic, and more harmonious music. Compared with MixLDM, we showed that modeling sources is much easier than directly modeling music mixtures. Compared with ISLDM, we showed that our model is able to model inter-source dependency or harmony between different musical instruments.
@@ V Conclusion and Future Work
In this paper, we propose a multi-source diffusion model to jointly model the generation of multiple instruments together. Both objective and subjective metrics show better results in the tasks of total and partial generation, implying MSLDM is able to efficiently model melody and inter-source relations. Future research needs to advance MSLDM for weakly-supervised music separation and generalize our model to more instruments.